DIFFERENTIAL GEOMETRY/PDE SEMINAR

Wednesday, January 22, 2003 Padelford C-36 3:45 pm

Exponential decay of eigenfunctions in many-body type scattering with second order perturbations

András Vasy (MIT)

In this talk I will discuss the exponential decay of eigenfunctions of secondorder geometric many-body type Hamiltonians at non-threshold energies, generalizing results of Froese and Herbst in potential scattering. I will also describe conditions that guarantee the absence of positive eigenvalues. The emphasis will be on the method, which is new in this setting and which has close ties to Hormander’s solvability condition for PDEs. The argument can be made lucid in the simple two-body setting, without losing the essential features that make it work in the many-body setting.

For more information about this seminar, visit the DG/PDE Seminar Web page (from the Math Department home page, www.math.washington.edu, follow the link Seminars, Colloquia, and Conferences).

The University of Washington is committed to providing access, equal opportunity and reasonable accommodation in its services, programs, activities, education and employment for individuals with disabilities. To request disability accommodation contact the Disability Services Office at least ten days in advance at: 206-543-6450/V, 206-543-6452/TTY, 206-685-7264 (FAX), or dso@u.washington.edu.